OpenStudy (anonymous):

If f(x) = 1 over 3x − 2 and g(x) = −2x − 1, find (f • g)(9)

OpenStudy (anonymous):

If r(x) = 3x − 1 and t(x) = −4x − 1, find (r + t)(7).

OpenStudy (anonymous):

If f(x) = −2x − 4 and g(x) = x + 4, find (f over g)(−8).

OpenStudy (anonymous):

I don't understand how to do them and it would be great if someone could show me how.